A material has conductivity of 10-2 mho/m and a relative permittivity of 4. The frequency at which conduction current in the medium is equal to displacement current is


1
Expert's answer
2021-01-14T10:41:12-0500

Let's first write the formula for the conduction current density:


"J=\sigma E,"

here, "J" is the conduction current density, "\sigma=10^{-2}\ \dfrac{S}{m}" is the conductivity of the material.

Then, we can write the formula for the displacement current density:


"J=\omega D=2\pi fD,"

here, "E" is the electric field, "\omega" is the angular frequency, "D=\epsilon\epsilon_0 E" is electric displacement field, "\epsilon=4" is a relative permittivity, "\epsilon_0=8.85\cdot10^{-12}\ \dfrac{F}{m}" is the permittivity of the free space, "f" is the frequency.

Since we searching for frequency at which conduction current in the medium is equal to displacement current, we must equate these two formulas:


"\sigma E=2\pi fE\epsilon\epsilon_0."

Finally, from this formula we can find the frequency at which conduction current in the medium is equal to displacement current:


"f=\dfrac{\sigma}{2\pi\epsilon},""f=\dfrac{10^{-2}\ \dfrac{S}{m}}{2\pi\cdot4\cdot8.85\cdot10^{-12}\ \dfrac{F}{m}}=45\ MHz."

Answer:

"f=45\ MHz."